At 25 degree Celsius, the vapour pressure of pure water is 23.76 mm Hg and that of an aqueous dilute solution of urea is 22.98 mm Hg .calculate the molality of this solution

Grade:12

2 Answers

varun
6 years ago
HELLO ABHISHEK,
we know that,
total vapour pressure = pure vapour pressure X mole fraction
22.98  =  23.76 X Xw
therefore,
Xw = 22.98/23.76
       =Nw/Nw+ Na
so,
22.98/23.76--22.98
====>Nw/Na
therefore,
Na/Nw = 23.76--22.98/22.98
 
NOW,
molality = Na/mass of water in kg
             =Na/Nw X 18X10^-3
 so molality = 23.76--22.98/22.98 X 18X10^-3
so the final answer = 0.001X 10^-3
THANK YOU
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
We know that,
total vapour pressure = pure vapour pressure X mole fraction
22.98  =  23.76 X Xw
therefore,
Xw = 22.98/23.76
       =Nw/Nw+ Na
so,
22.98/23.76--22.98
====>Nw/Na
therefore,
Na/Nw = 23.76--22.98/22.98
 
NOW,
molality = Na/mass of water in kg
             =Na/Nw X 18X10^-3
 so molality = 23.76--22.98/22.98 X 18X10^-3
so the final answer = 0.001X 10^-3
 
Hope it helps.
Thanks and regards,
Kushagra